# FFGAF-SNN: The Forward-Forward Based Gradient Approximation Free Training Framework for Spiking Neural Networks

### v1

@@ Introduction
Applying FF algorithms to SNN training offers an effective alternative to traditional backpropagation. Contrastive Signal-Dependent Plasticity (CSDP)(Ororbia 2023) scheme based on FF learning rules, introducing activation trajectories for SNN self-supervised learning, boosting performance and enabling parallel learning/inference execution. Terres designed an SNN training framework based on the Forward-Forward algorithm for robust out-of-distribution detection, formulating SNN-compatible “goodness” functions that achieve energy-efficient OoD detection(Terres-Escudero et al. 2024). Backpropagation-free Spiking Neural Networks with the Forward-Forward Algorithm(Ghader et al. 2025) developed a FF algorithm-based SNN training framework using two forward passes for inter-layer local learning, improving computational efficiency and hardware compatibility while delivering superior performance across multiple datasets. In summary, FF algorithms enhance SNN training efficiency, biological plausibility, and energy efficiency, showing great application potential. However, existing methods still fail to resolve the non-differentiability of neuronal activation functions and have relatively low implementation accuracy.
@@
$\tilde{s}_{c}=\frac{\sum_{c=1}^{K}\mathbf{S}_{c,c^{\prime}}-\mu_{s}}{\sigma_{s}}$ (3)
@@
$S_{j}=\left\lfloor\frac{\tilde{s}c-\min{\tilde{s}_{c}}+\phi}{\sum_{c=1}^{C}(\tilde{s}_{c}-\min{\tilde{s}_{c}}+\phi)}\cdot C_{total}\right\rfloor$ (4)
@@
To evaluate the performance of convolutional layers, we assume there are $J$ data classes. The feature map activation outputs of each convolutional layer are denoted as $\mathbf{Y}\in\mathbb{R}^{N\times C\times H\times W}$, where $N$ is the number of samples in the mini-batch, $C$ is the channel dimension, and $H$, $W$ are the spatial dimensions. We refer to (Papachristodoulou et al. 2024), assuming $J$ data classes, $\mathbf{Y}$ is partitioned into $J$ subsets, with the number of channels in each subset determined by the complexity of the corresponding class $j$. Thus, $\hat{\mathbf{Y}}_{j}\in\mathbb{R}^{N\times S_{j}\times H\times W}$ represents the activation output of $j^{th}$ class, where $S_{j}$ is the number of channels allocated to class $j$. The holistic goodness factor for each layer is a matrix $\mathbf{G}\in\mathbb{R}^{N\times J}$. Each element $G_{n,j}$ in this matrix quantifies the goodness of class $j$ for example $n$, computed via the spatial and channel-wise mean of the square of $\hat{\mathbf{Y}}_{j}$, as shown below.
$G_{n,j}=\frac{1}{S_{j}\times H\times W}\sum_{s=1}^{S_{j}}\sum_{h=1}^{H}\sum_{w=1}^{W}\hat{Y}_{n,j,s,h,w}^{2}$ (5)
@@
$\text{Regularized Output}=\text{thresh}\times\frac{\text{Output}-\mu}{\sigma}$ (10)
@@
This section details the experiments conducted to evaluate the proposed techniques and compare them with existing research. Model performance is assessed across three benchmark datasets commonly used in literature: MNIST (LeCun, Cortes, and Burges 1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), and CIFAR-10 (Krizhevsky and Hinton 2009). The model was trained on an NVIDIA-TESLA V100 GPU for 50 epochs on MNIST and Fashion-MNIST, and for 100 epochs on CIFAR-10. During the experiments, the learning rate for the convolutional layers was set to 0.01, with a batch size of 128. The convolutional stride was configured as [1, 2, 1, 2, 1], and the network architecture comprised layers with dimensions [40, 120, 120, 240].
@@
Table 1: This table describes the test accuracy and time steps of different training algorithms and network types across various datasets. FF* and CwComp* represent the accuracy achieved by our implemented models.
@@
We present comprehensive experimental results comparing our proposed Forward-Forward Gradient Approximation Free Training Framework (FFGAF-SNN)with state-of-the-art approaches. The FFGAF-SNN architecture employs an innovative design comprising one encoding layer training block and three hidden layer training blocks, each followed by an SNN IF neuron layer for output encoding and activation. To address feature information loss in the SNN neuron layer, we developed an enhanced ReLU function and incorporated channel-wise weight parameters for improved feature discrimination. The architecture utilizes batch dimension normalization in the first training block and temporal dimension regularization in hidden layers, demonstrating superior performance across MNIST, Fashion-MNIST, and CIFAR-10 datasets while significantly reducing model parameters. Notably, our framework achieves particularly remarkable results on CIFAR-10. It represents the first successful implementation of forward-forward training on CSNNs, offering substantial reductions in both parameter count and memory accesses compared to conventional FF-trained SNNs. The experimental results show that our method outperforms existing non-backpropagation learning algorithms, including those based on Direct Feedback Alignment (DFA-based) and Spike-Timing-Dependent Plasticity (STDP-based) approaches. Furthermore, our algorithmic model requires fewer iterations and converges faster than most models. Our model not only surpasses ANN models trained with DFA (FFF, DRTP), but also significantly exceeds the accuracy of Spiking Neural Network models trained with DFA (aDFA) by nearly 5% on the Fashion-MNIST dataset. Compared with STDP-based SNN training methods (sym-STDP, SDNN), our approach also demonstrates higher accuracy. Testing on the Fashion-MIST dataset indicates that our model outperforms sym-STDP by nearly 7%. Additionally, comparative studies with models trained using backpropagation show that our FFGAF-trained SNN can achieve comparable accuracy while eliminating the need for backpropagation (3% higher on the Fashion-MINST dataset).
@@
A detailed energy efficiency analysis highlights the practical advantages of our FFGAF-SNN model. Compared to MLP networks trained with FF, this architecture maintains a significantly smaller parameter footprint—only 2.3% of the parameter count of FF* models—while achieving approximately a 20% improvement in accuracy on the CIFAR-10 dataset. Thanks to its fewer parameters, memory access energy consumption is substantially reduced relative to SNN models trained with FF. The memory access energy is only 14% of the model proposed by Ororbia, and it remains favorable when compared to other CNN architectures. Although computational energy benefits from the inherent efficiency of SNN operations, the convolutional implementation results in slightly higher inference energy consumption than some SNN alternatives. However, compared to CNN models, our model’s computational energy is only 9% of CwComp* and our model’s total energy consumption of 0.081mJ is significantly lower than other SNN models, which consume 0.175mJ and 0.325mJ. These combined advantages position our model as an exceptionally efficient solution for edge deployment scenarios.

### v2

@@ Introduction
Applying FF algorithms to SNN training offers an effective alternative to traditional backpropagation. Contrastive Signal-Dependent Plasticity (CSDP)(Ororbia 2024) scheme based on FF learning rules, introducing activation trajectories for SNN self-supervised learning, boosting performance and enabling parallel learning/inference execution. Terres designed an SNN training framework based on the Forward-Forward algorithm for robust out-of-distribution detection, formulating SNN-compatible “goodness” functions that achieve energy-efficient OoD detection(Terres-Escudero et al. 2024). Backpropagation-free Spiking Neural Networks with the Forward-Forward Algorithm(Ghader et al. 2025) developed a FF algorithm-based SNN training framework using two forward passes for inter-layer local learning, improving computational efficiency and hardware compatibility while delivering superior performance across multiple datasets. In summary, FF algorithms enhance SNN training efficiency, biological plausibility, and energy efficiency, showing great application potential. However, existing methods still fail to resolve the non-differentiability of neuronal activation functions and have relatively low implementation accuracy.
@@
$\tilde{s}_{c}=\frac{\sum_{c^{\prime}=1}^{K}\mathbf{S}_{c,c^{\prime}}-\mu_{s}}{\sigma_{s}}$ (3)
@@
$S_{j}=\left\lfloor\frac{\tilde{s}_{j}-\min{\tilde{s}_{j}}+\phi}{\sum_{c=1}^{K}(\tilde{s}_{c}-\min{\tilde{s}_{c}}+\phi)}\cdot C_{total}\right\rfloor$ (4)
@@
[…]
$G_{n,j}=\frac{1}{S_{j}\times H\times W}\sum_{s=1}^{S_{j}}\sum_{t=1}^{T}\sum_{h=1}^{H}\sum_{w=1}^{W}\hat{Y}_{n,j,t,s,h,w}^{2}$ (5)
@@
[…]
@@
This section details the experiments conducted to evaluate the proposed techniques and compare them with existing research. Model performance is assessed across three benchmark datasets commonly used in literature: MNIST (LeCun, Cortes, and Burges 1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), and CIFAR-10 (Krizhevsky and Hinton 2009). The model was trained on an NVIDIA-TESLA V100 GPU for 30 epochs on MNIST,50 epochs on Fashion-MNIST, and for 100 epochs on CIFAR-10. The time step is configured to be 10.During the experiments, the learning rate for the convolutional layers was set to 0.01, with a batch size of 128. The convolutional stride was configured as [1, 2, 1, 2, 1], and the network architecture comprised layers with dimensions [40, 120, 120, 240].
@@
Table 1: This table describes the test accuracy and training epochs of different training algorithms and network types across various datasets. FF* and CwComp* represent the accuracy achieved by our implemented models.
@@
We present comprehensive experimental results comparing our proposed Forward-Forward Gradient Approximation Free Training Framework (FFGAF-SNN)with state-of-the-art approaches. The FFGAF-SNN architecture employs an innovative design comprising one encoding layer training block and three hidden layer training blocks, each followed by an SNN IF neuron layer for output encoding and activation. To address feature information loss in the SNN neuron layer, we developed an enhanced ReLU function and incorporated channel-wise weight parameters for improved feature discrimination. The architecture utilizes batch dimension normalization in the first training block and temporal dimension regularization in hidden layers, demonstrating superior performance across MNIST, Fashion-MNIST, and CIFAR-10 datasets while significantly reducing model parameters. Notably, our framework achieves particularly remarkable results on CIFAR-10. It represents the first successful implementation of forward-forward training on CSNNs, offering substantial reductions in both parameter count and memory accesses compared to conventional FF-trained SNNs. As shown in Table 1, the experimental results show that our method outperforms existing non-backpropagation learning algorithms, including those based on Direct Feedback Alignment (DFA-based) and Spike-Timing-Dependent Plasticity (STDP-based) approaches. Furthermore, our algorithmic model requires fewer iterations and converges faster than most models. Our model not only surpasses ANN models trained with DFA (FFF, DRTP), but also significantly exceeds the accuracy of Spiking Neural Network models trained with DFA (aDFA) by nearly 5% on the Fashion-MNIST dataset. Compared with STDP-based SNN training methods (sym-STDP, SDNN), our approach also demonstrates higher accuracy. Testing on the Fashion-MIST dataset indicates that our model outperforms sym-STDP by nearly 7%. Additionally, comparative studies with models trained using backpropagation show that our FFGAF-trained SNN can achieve comparable accuracy while eliminating the need for backpropagation (3% higher on the Fashion-MINST dataset). Additionally, our algorithm also outperformed FF-based ANN models such as PFF(Ororbia and Mali 2023), FFCL(Karkehabadi, Homayoun, and Sasan 2024), CaFo(Zhang et al. 2024), and so on, demonstrating the superiority of our approach.
@@
As shown in Table 2, a detailed energy efficiency analysis highlights the practical advantages of our FFGAF-SNN model. Compared to MLP networks trained with FF, this architecture maintains a significantly smaller parameter footprint—only 2.3% of the parameter count of FF* models—while achieving approximately a 20% improvement in accuracy on the CIFAR-10 dataset. Thanks to its fewer parameters, memory access energy consumption is substantially reduced relative to SNN models trained with FF. The memory access energy is only 14% of the model proposed by Ororbia, and it remains favorable when compared to other CNN architectures. Although computational energy benefits from the inherent efficiency of SNN operations, the convolutional implementation results in slightly higher inference energy consumption than some SNN alternatives. However, compared to CNN models, our model’s computational energy is only 9% of CwComp* and our model’s total energy consumption of 0.081mJ is significantly lower than other SNN models, which consume 0.175mJ and 0.325mJ. These combined advantages position our model as an exceptionally efficient solution for edge deployment scenarios.